OpenStudy (anonymous):

(3m^2-2m-35)/(m^2-6m+9) + (m+4)/(m-7) Find the LCD

OpenStudy (austinl):

\[\frac{(3m^2-2m-35)}{(m^2-6m+9)} + \frac{(m+4)}{(m-7)}\] You would need to multiply the denominators together.

OpenStudy (austinl):

\[(m^2-6m+9)\times (m-7)\]

OpenStudy (anonymous):

So m^3-13m^2+51m-63???

OpenStudy (anonymous):

Do you multiply the denominators and then factor them??

OpenStudy (austinl):

I made an error.... \[\frac{m-7}{m-7}\times \frac{(3m2−2m−35)}{(m2−6m+9)}+\frac{(m+4)}{(m−7)}\times \frac{m^2-6m+9}{m^2-6m+9}\] I believe this is what you need to do, apologies, I need caffeine.

OpenStudy (anonymous):

Why would I do that? I'm confused now. XD

OpenStudy (austinl):

Perhaps it'd be best if we got an expert in here. I am too tired to be doing this. I'm sorry.

OpenStudy (austinl):

@nincompoop care to take over?

OpenStudy (anonymous):

It is okay.

hero (hero):

Try factoring first before adding